40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 21 02:42:17 2014, rana, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00002.PNGTEK00003.PNG
    Reply  Thu Aug 21 19:04:52 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00005.PNG
       Reply  Tue Aug 26 23:16:49 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH Ytfs.pdfpdhComparison.pdfnewY.zip
Message ID: 10420     Entry time: Thu Aug 21 19:04:52 2014     In reply to: 10418     Reply to this: 10430
Author: ericq 
Type: Configuration 
Category: Green Locking 
Subject: Gain changes on Green Y PDH 

I found that the barrel of one the BNC to BNC connectors used for getting the output of the PDH servo box to the laser controller was touching the ETMY chamber. When I held it away, all of the 60Hz harmonics disappeared from the mixer output spectrum; this was pretty repeatable. This inspired me to replace the refl PD and PZT signal cables (which were 2 and 3 cables stitched together, respectively) with 20' long BNCs. I also cleaned up a lot of the routing of signal and power cables in the little rack, and moved the big T->DC Block->Attenuator combo off of the panel mount, because I didn't like how it was wiggling. It and the summing pomona box are sitting on top of the PDH box and function generator, instead of hanging freely.

All of the 60Hz harmonics were banished afterwards, and the green locked happily. 

This required me touching the Y end table, to remove the old cable and its cable ties, and putting the new one in. I don't think I did anything immediately apparently bad; the green and IR transmissions both are within nominal ranges. 

I haven't had luck measuring the CLG yet, which I wanted to do to get and set the UGF before measuring the noises. However, here is a scope trace of the in-lock error signal, which compares quite favorably to the trace posted in the previous post; the scope indicates that the signal has 1/3 of the RMS that it did before I replaced the cables. 

TEK00005.PNG

I hope to measure up the current status after I get back from dinner. 

 

ELOG V3.1.3-